A student decides to perform the synthesis of triphenylmethanol via the grignard reaction. She starts with 9.5g of bromobenzene (M.W. = 157) and uses 1.21g of Mg (M.W. = 24.3) and 9.1g of benzophenone (M.W. = 182).

a. Calculate the limiting reagent.

b. Calculate the theoretical yield of triphenylmethanol (M.W. = 260.3).
Answer #1

Solution:

A) Calculation of limiting reagent:

Number of moles of bromobenzene = Mass /Molar mass

= 9.5 g / 157 g mol-1 = 0.0605 mol

Number of moles of Mg = Mass /Atomic mass

= 1.21 g / 24.3 g mol-1 = 0.0498 mol

Number of moles of benzophenone = Mass / Molar mass

= 9.1 g / 182 g mol-1 = 0.050 mol

Since, number of moles of Mg is minimum, hence Mg is the limiting reagent.

B) Theoretical yield of triphenylmethanol:

Since, overall synthesis is 1:1 stoichiometric, hence 0.0498 mol of Mg will form 0.0498 mol of triphenylmethanol.

Therefore,

Theoretical yield = Number of mol x molar mass

= 0.0498 mol x 260.3 g mol-1 = 12.963 g